OpenStudy (anonymous):

Write the equation of the line parallel to the line 4x + y = -1 and passing through the point (5,0)

OpenStudy (anonymous):

For this problem, use this 4x + y = ? Plug in the point (5, 0) and solve. 4(5) + 0 = ? What you get for the question mark, put it where the blank is down here: 4x + y = __

OpenStudy (anonymous):

20

OpenStudy (anonymous):

Yup. The answer is 4x + y = 20

OpenStudy (anonymous):

So then what would you do for a problem like this? Write the equation of the line parallel to the line 10x - 5y = 8 and passing through the point (2,4).

OpenStudy (anonymous):

Same thing. If it says PARALLEL and it's in the form Ax + By = C, then do the same thing.